OpenStudy (daniel.ohearn1):

Find the point on the hyperbola xy=10 closest to the point (3,0). Is there a way to do this without looking at a graph?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

probably calculus

OpenStudy (daniel.ohearn1):

I hope so, it's for my Calculus 1A class..

OpenStudy (misty1212):

lol then that is what you can use

OpenStudy (misty1212):

a point on the graph will look like \((x,\frac{10}{x})\)

OpenStudy (misty1212):

and the square of the distance between that point and \((3,0)\) will be \[\large (x-3)^2+\left(\frac{10}{x}\right)^2\] minimize that in the usual calc way

OpenStudy (daniel.ohearn1):

I know I should take a derivative right?

OpenStudy (daniel.ohearn1):

implicitly or otherwise with respect to x?

OpenStudy (daniel.ohearn1):

If I take a derivative and solve it for zero I get 2x^4 - 6x^3 = 0 .... hmm..

OpenStudy (daniel.ohearn1):

I got x=(1/3)^(5/4) but that can't be right.. I'm not sure how to do this atm..

OpenStudy (daniel.ohearn1):

Help?